6 2/9 divided by 5 1/9 - 7/8 = ?
Answer by edjones(8007)

6 2/9 divided by (5 1/9 - 7/8) I assume that this is the problem.
=56/9 divided by (46/9 - 7/8)
=56/9 d b (368/72 - 63/72)
=56/9 d b 305/72
=56/9 * 72/305
=56*8/305
=448/305
.
Ed